# Predicting sampling advantage of stochastic Ising Machines for Quantum Simulations

### v1

@@ III Discussion
To better understand the increased autocorrelation at $\alpha>2$ as compared to $\alpha=2$, we studied the differences between the probabilistic flipping of visible and hidden spins. This is of key relevance since we are considering the variational quantum energy as key observable, which depends only on visible spins. The increased autocorrelation time, which is reflected in the freezing of the visible spins, can be explained by the height of the energy barrier required to flip a visible spin. The energy barrier is defined as $\Delta{E_{i}}=H_{\text{ising}}(m_{i}\rightarrow-m_{i})-H_{\text{ising}}(m)$ and the height is determined by the coupling strength of a visible spin to all the $M=\alpha n$ hidden spins. In supplement E, we show that the mean coupling strength decreases at a slower rate than $\alpha$ increases. Combined with the fact that connectivity per visible spin increases with $\alpha$, gives a net growth of the energy barrier and therefore exponential suppression of flipping visible spins. We further confirm this interpretation by observing a significantly higher rate of probabilistic flipping of the hidden spins, consistent with their connectivity remaining constant as $\alpha$ increases. Hence, the hidden and visible spins feature very much distinct dynamics. This explanation suggests great potential for using sparse DBMs, which may suffer much less from the high energy barriers, while still offering great expressiveness [36].
@@
In addition to the runtime estimates, we can also make a back of the envelope energy comparison. For the conservative projection, we use the largest system size ($n_{\mathrm{spins}}=484$, $\alpha=2$) and that the energy usage per AIMC core is $100$ mW [33], 4 cores are used and the device requires 400 ns per step. For MH (UltraFast) on an AMD EPYC 7601 CPU, we use the reported 180 W TDP divided by the 32 cores and measured average runtime per core while utilizing 16 cores. With these settings, the conservative projection is estimated to be 2800 times more energy efficient than MH (UltraFast) on CPU.
@@
Given the highly promising sampling advantages, we expect that our studies open up new opportunities for large scale quantum simulations. Particularly interesting next steps include investigation of sampling advantages far away from the ground states, which is relevant for training NQS. Further changing the topology of the Ising model to models such as sparse DBMs, could possibly be used to tune the autocorrelation time without sacrificing accuracy. Another direction is to study frustrated models as well as unitary quantum dynamics on the sIM, which would involve using an ansatz with amplitude and phase.
@@ B.1 Training parameters
Metropolis-Hastings (MH) sampling was used for computing expectation values. The thermalization time was configured to $200$ sweeps and samples were taken one sweep apart, where the sweep size is equal to the number of visible spins. The number of training iterations was between $300$ and $600$ iterations and the learning rate $\eta=0.005$. For the sampling settings and regularization two different sets of parameters were used depending on the number of spins and $\alpha$, which is given in Table 1. There is a set with ’low’ which corresponds to $2000$ MC samples and ’high’ corresponds to $10000$ MC samples. Also the regularization is different for these two sets and is given by:
@@
Table 1: Summary of hyperparameters depending on $n_{\mathrm{spins}}$ and $\alpha$ for training with standard and modified RBM ansatze.
@@
B.2 Accuracy of modified RBM
The accuracy of the modified RBM ansatz, Eq. (3), compared to the ’standard’ RBM ansatz used in [9, 10] is presented in Fig. 5. The accuracy is expressed using the relative error given by
$\varepsilon_{\mathrm{rel}}=\left|\frac{E_{\mathrm{NQS}}-E_{\mathrm{QMC}}}{E_{\mathrm{QMC}}}\right|$ (21)
@@ B.2 Accuracy of modified RBM
Figure 5: The relative error, Eq. (21), as a function of the side length $L$ of 2D antiferromagnetic Heisenberg model. The QMC ground state energies are taken from [40] for comparison. Standard refers to the ansatz used in [9, 10]. The model id’s, 00 through $4$, correspond to the 5 pre-trained models.
For all the models, $32$ chains of $10000$ MH sweeps ($n_{\mathrm{spins}}$ times updating two spins) were used. For some of these models the Markov chain got stuck, these chains were excluded from the comparison and are listed in Table 2.
@@ Appendix C Autocorrelation time
The autocorrelation time of the variational energy was measured for the regular MH and sIM sampling. To measure the autocorrelation time, the autocovariance function needs to be calculated which is defined as
$\Gamma_{c}=\frac{1}{N_{\mathrm{S}}-c}\sum_{i=1}^{N_{\mathrm{S}}-c}(O_{i}-\overline{O})(O_{i+c}-\overline{O})$ (22)
where $N_{\mathrm{S}}$ is the number of samples in the Markov chain and $c$ is the lag time. The autocorrelation function is the normalized autocovariance function and is defined as
$\rho_{c}=\frac{\Gamma_{c}}{\Gamma_{0}}.$ (23)
Note that $\Gamma_{0}$ is equal to the variance of the observable $O$ [41].
The autocorrelation time can be approximated by integrating the autocorrelation function like
$\tau_{\mathrm{int}}=\frac{1}{2}+\sum_{c=1}^{M}\rho_{c}$ (24)
where $M$ is a cut-off value for the integration. When $c$ starts to become closer to the length of the Markov chain $N_{\mathrm{S}}$, the autocorrelation function $\rho$, Eq. (23), becomes inaccurate and small and therefore the sum can be truncated, given that there are at least a couple of autocorrelation times worth of data in the chain. The cutoff value is the smallest $M$ such that $M\geq 4\tau_{\mathrm{int}}+1$ and for that $M$ we have $\tau=\tau_{\mathrm{int}}$. This approach is taken from [42].
The following sampling settings were used for calculating the autocorrelation time. For the MH sampling, the sampling interval (sweep size) was set to
$\mathrm{sweep\_size}=\max(0.01\cdot n_{\mathrm{spins}},1)$ (25)
@@
For sIM, the sampling interval was taken at a variable interval. The interval was chosen such that a chain contained at least $1500$ magnetization 0 autocorrelation times and the autocorrelation time measured in the samples was at least bigger than $2$ (except for $\alpha\leq 2$ where at least bigger than $1.5$). There are at least $5$ chains of at least $32768$ samples taken for each model. The autocorrelation time for both MH and sIM was only calculated on states within the magnetization sector 00. In Fig. 6 the autocorrelation time of sIM is shown expressed in sIM sweeps, where a sIM sweep corresponds to one update of all visible spins and one update of all hidden spins. In Fig. 7, the autocorrelation time of the MH sampling is shown expressed in sweeps, which corresponds to $n_{\mathrm{spins}}$ times updating two spins. Some models suffered from Markov chains getting stuck, which are listed in the sIM and MH columns of Table 2. Additionally, results for $\alpha=1$ are already excluded due to poor performance.
Figure 6: Autocorrelation time of the variational energy for sIM sampling. For every combination of $\alpha$ and $n_{\mathrm{spins}}$ there is averaging over $5$ pre-trained models.
Figure 7: Autocorrelation time of the variational energy for the Metropolis-Hastings sampling. For every combination of $\alpha$ and $n_{\mathrm{spins}}$ there is averaging over $5$ pre-trained models.
Table 2: Models excluded from analysis due to Markov chain getting stuck for either sIM or MH sampling. $\alpha=1$ has been excluded entirely due to poor performance.
$n_{\mathrm{spins}}$ $\alpha$ Id sIM MH
64 1 0 x
64 1 1 x x
100 1 0 x
100 1 2 x
100 1 3 x
100 1 4 x
@@
Appendix D Performance sIM GPU
@@
The performance of a single chain sIM (chromatic Gibbs) GPU implementation written in PyTorch on an NVIDIA L40S GPU was benchmarked for random weights and biases. In Fig. 8, the performance of that GPU implementation is compared to the single chain sIM (chromatic Gibbs) CPU implementation written in Julia. The GPU implementation is only competitive for larger system sizes ($n_{\mathrm{spins}}>700$) for $\alpha=2$.
@@
Figure 8: Comparison of the time per step for a single chain sIM GPU implementation and a single chain sIM CPU implementation as a function of the number of visible spins $n_{\mathrm{spins}}$ for $\alpha=2$. Both implementations use chromatic Gibbs sampling and a step refers to one update of all visible spins and one update of all hidden spins. The GPU implementation is written in PyTorch and the CPU implementation is written in Julia.
Appendix E Energy barrier analysis
@@
$\Delta E_{i}(m)=H_{\text{ising}}(m_{i}\rightarrow-m_{i})-H_{\text{ising}}(m)$ (26)
where $E(m)$ is the energy of the state $m$ given by Eq. (4). The high autocorrelation time for models with higher $\alpha$ is related to the freezing of the visible spins which suggests that there is a high energy barrier for flipping the visible spins. In Fig. 9a, the average measured energy barrier for flipping is shown as a function of $n_{\mathrm{spins}}$ and $\alpha$. The energy barrier can be compared to the logarithm of the autocorrelation time shown in Fig. 9b, which shows qualitatively the same behavior. This suggests that the autocorrelation time is related to the energy barrier of the visible spins. Additionally, the energy barrier of the visible spins can be roughly approximated. For that we first simplify Eq. (26) to
$\Delta E_{i}(m)=2m_{i}\sum_{j=1}^{M}J_{ij}m_{j}+h_{i}m_{i}$ (27)
@@
Figure 9: (a) The average energy barrier associated with flipping visible spins, given by $\Delta E=\frac{1}{n}\sum_{i=1}^{n}\frac{1}{N}\sum_{\mu=1}^{N}\Delta E_{i}(m^{\mu})=\frac{1}{nN}\sum_{\mu=1}^{N}\sum_{i=1}^{n}E(m_{i}^{\mu}\rightarrow-m_{i}^{\mu})-E(m^{\mu})$. (b) The logarithm of the autocorrelation time $\tau_{\mathrm{Evar,sIM}}$. (c) The approximated energy barrier over the weight matrix ${\Delta E^{\mathrm{approx}}_{s_{i}}}=\frac{1}{n}\sum_{i=1}^{n}\sum_{j=1}^{M}|W_{ij}|$ as a function of $n_{\mathrm{spins}}$ and $\alpha$. (a, b, c) The models considered have Id = 0 and the results are presented as a function of $\alpha$ and $n_{\mathrm{spins}}$.
Figure 10: The mean connection strength of the weight matrix, given by $\frac{1}{\alpha n^{2}}\sum_{i,j=1}^{n,M}|W_{ij}|$ as a function of $n_{\mathrm{spins}}$ and $\alpha$

### v2

@@ III Discussion
To better understand the increased autocorrelation at $\alpha>2$ as compared to $\alpha=2$, we studied the differences between the probabilistic flipping of visible and hidden spins. This is of key relevance since we are considering the variational quantum energy as key observable, which depends only on visible spins. The increased autocorrelation time, which is reflected in the freezing of the visible spins, can be explained by the height of the energy barrier required to flip a visible spin. The energy barrier is defined as $\Delta{E_{i}}=H_{\text{ising}}({z}_{i}\rightarrow-{z}_{i})-H_{\text{ising}}({z})$ and the height is determined by the coupling strength of a visible spin to all the ${m}=\alpha n$ hidden spins. In Appendix H, we show that the mean coupling strength decreases at a slower rate than $\alpha$ increases. Combined with the fact that connectivity per visible spin increases with $\alpha$, gives a net growth of the energy barrier and therefore exponential suppression of flipping visible spins. We further confirm this interpretation by observing a significantly higher rate of probabilistic flipping of the hidden spins, consistent with their connectivity remaining constant as $\alpha$ increases. Hence, the hidden and visible spins feature very much distinct dynamics. This explanation suggests great potential for using sparse DBMs, which may suffer much less from the high energy barriers, while still offering great expressiveness [39].
@@
In addition to the runtime estimates, we can also make a back of the envelope energy comparison. For the conservative projection, we use the largest system size (n=484{n}=484, $\alpha=2$) and that the energy usage per AIMC core is $100$ mW [34], 4 cores are used and the device requires 400 ns per step. For MH (UltraFast) on an AMD EPYC 7601 CPU, we use the reported $180$ W TDP divided by the $32$ cores and measured average runtime per core while utilizing $16$ cores. With these settings, the conservative projection yields about $3$ orders of magnitude higher energy efficiency than MH (UltraFast) on CPU. The TDP gives a rough estimate of the energy usage, in future work, more accurate energy comparisons can be obtained by directly measuring the energy, which is possible using the energy aware runtime (EAR) [40].
@@
Given the highly promising sampling advantages, we expect that our studies open new opportunities for large scale quantum simulations. Particularly interesting next steps include investigation of sampling advantages far away from the ground states, which is relevant for training NQS. Further changing the topology of the Ising model to models such as sparse DBMs [39] is very interesting. DBMs have the same blocking structure as RBMs and hence our methodology can straightforwardly be applied to such models as well and potentially the sparsity can be used to tune the autocorrelation time without sacrificing accuracy. Another direction is to study frustrated models as well as unitary quantum dynamics on the sIM, which would involve using an ansatz with amplitude and phase. Beyond applications to NQS, the methodology developed in this paper can straightforwardly be used for estimating possible speedups of hardware improvements on Boltzmann machine simulation of general machine learning workloads, which is particularly interesting in view of the rapid development of Ising machine hardware.
@@ B.1 Training parameters
Metropolis-Hastings (MH) sampling was used for computing expectation values. The thermalization time was configured to $200$ MH sweeps and samples were taken one sweep apart. The number of training iterations was between $300$ and $600$ iterations and the learning rate $\eta=0.005$. For the sampling settings and regularization two different sets of parameters were used depending on the number of spins and $\alpha$, which is given in Table 1. There is a set with ’low’ which corresponds to $2000$ MC samples and ’high’ corresponds to $10000$ MC samples. Also the regularization is different for these two sets and is given by:
@@
Table 1: Summary of hyperparameters depending on n{n} and $\alpha$ for training with standard and modified RBM ansatze.
@@
B.2 Accuracy of pre-trained models
The accuracy of the ground state energy of the pre-trained RBM ansatz is compared to the QMC ground state energy from [31] in Fig. 5. The accuracy is expressed using the relative error given by
$\eta_{\mathrm{rel}}=\left|\frac{E_{\mathrm{NQS}}-E_{\mathrm{QMC}}}{E_{\mathrm{QMC}}}\right|$ (21)
@@ B.2 Accuracy of pre-trained models
Furthermore, the accuracy of the pre-trained models with the wave function ansatz given in Eq. (3), are compared to the ’standard’ RBM ansatz used in [8, 9] in Fig. 5. For all the models, $32$ chains of $10000$ MH sweeps were used. For some of these models the Markov chain got stuck in samples with high $E_{\mathrm{loc}}$, we excluded these chains based on the average quantum energy: E>0.55E>-0.55, which is much higher than results from e.g. Linear Spin Wave Theory E=0.658E=-0.658 per spin [44]. The relative error was then calculated over the remaining chains. The models where chains are excluded are listed in the MH columns of Table 2.
Figure 5: The relative error, Eq. (21), as a function of the side length $L$ of 2D antiferromagnetic Heisenberg model. The QMC ground state energies are taken from [31] for comparison. Colored symbols refer to the results obtained from 5 different pretrained models per $\alpha$ and n{n}. Black circles with error bars are results obtained with ansatz used in [8, 9].
@@ Appendix C Autocorrelation time
Using Eq. (14), we can estimate the sampling interval from the autocorrelation time and therefore the number of steps required on a sIM to reach iso-accuracy compared to MH sampling. To find the autocorrelation time from the autocorrelation function $\rho(j)$, we use the integrated autocorrelation time approach, where the autocorrelation function is integrated up to a cut-off value $C$:
$\tau_{\mathrm{int}}=\frac{1}{2}+\sum_{j=1}^{C}\rho(j).$ (22)
The cutoff value is the smallest $C$ such that $C\geq 4\tau_{\mathrm{int}}+1$ and for that $C$ we have $\tau\approx\tau_{\mathrm{int}}$ [45]. In Fig. 6 (7) the measured autocorrelation time of sIM (MH) is shown expressed in sIM (MH) sweeps respectively.
The following sampling settings were used for calculating the autocorrelation time. For the MH sampling, the sampling interval was set to
${\Delta t}=\max(0.01\cdot{n},1)\ \text{MH flips}$ (23)
@@
For sIM, the sampling interval was taken at a variable interval. The interval was chosen such that a chain contained at least $1500$ autocorrelation times and the autocorrelation time measured in the samples was at least larger than $2$ (except for $\alpha\leq 2$ where at least larger than $1.5$). There are at least $5$ chains of at least $32768$ samples taken for each model.
As mentioned in Appendix B.2, some models suffered from Markov chains getting stuck. This also occured for measuring the autocorrelation time of MH and sIM sampling. Here, instead of excluding chains, we excluded the entire pre-trained model if any of the chains had an average energy E>0.55E>-0.55 for either sIM and/or MH sampling. The excluded models are listed in Table 2. Additionally, results for $\alpha=1$ are already excluded due to poor performance.
Figure 6: Autocorrelation time of the variational energy for sIM sampling. For every combination of $\alpha$ and n{n} there is averaging over 55 pre-trained models.
Figure 7: Autocorrelation time of the variational energy for the Metropolis-Hastings sampling. For every combination of $\alpha$ and n{n} there is averaging over 55 pre-trained models.
Figure 8: The ratio of samples in the magnetization 0 sector $f$ sampled by sIM sampling as a function of the number of spins $n$ and side length $L$ for different hidden layer densities $\alpha$.
Table 2: Models excluded from the analysis. Model Id=0…4 refer to labels used in the dataset for 5 different models per $n$ and $\alpha$. The criterion for exclusion is based on the error in the relative energy, see the discussion in Appendix C
n{n} $\alpha$ Id sIM MH
@@
Appendix D Blocking Gibbs sampling
In this article, we implement blocking Gibbs sampling which allows to update all the hidden spins in parallel based on all the visible spins and vice versa. The blocking Gibbs sampling can be implemented on hardware Ising Machines and we have implemented it in Julia for the purpose of this article [35].
The blocking Gibbs sampling starts with a random initialization of the visible spins $s$ and then iteratively updates the hidden spins $x$ and visible spins $s$. The hidden spins are updated based on the input from the visible spins and is given by the following equation:
$I_{j}$ $=\sum_{i=1}^{n}W_{ij}s_{i}+b_{j}$ (24)
$x_{j}$ $=\mathrm{sgn}\left(\tanh\left(I_{j}\right)-r_{j}\right)$ (25)
where $r_{j}$ are $M$ uniform random numbers between 1-1 and $1$. The visible spins are updated based on the input from the hidden spins and is given by the following equation:
$I_{i}$ $=\sum_{j=1}^{M}W_{ij}x_{j}$ (26)
$s_{i}$ $=\mathrm{sgn}\left(\tanh\left(I_{i}\right)-r_{i}\right)$ (27)
where $r_{i}$ are $n$ uniform random numbers between 1-1 and $1$.
Appendix E Performance sIM GPU
@@
The performance of a single chain sIM (chromatic Gibbs) GPU implementation written in PyTorch on an NVIDIA L40S GPU was benchmarked for random weights and biases [35]. In Fig. 9, the performance of that GPU implementation is compared to the single chain sIM (chromatic Gibbs) CPU implementation written in Julia. The GPU implementation is only competitive for larger system sizes (n>700{n}>700) for $\alpha=2$.
@@
Figure 9: Comparison of the time per sIM update for a single chain sIM GPU implementation and a single chain sIM CPU implementation as a function of the number of visible spins n{n} for $\alpha=2$. Both implementations use chromatic Gibbs sampling and a sIM update refers to one update of all visible spins and one update of all hidden spins. The GPU implementation is written in PyTorch and the CPU implementation is written in Julia.
Appendix F Estimating the performance of the (hardware) sIM
Implementing blocking Gibbs sampling using stochastic Magnetic Tunnel Junctions (MTJs) implements Eq. 25 and 27 with stochastic MTJs. The matrix multiplication in Eq. 24 and 26 can be implemented using a crossbar. For the conservative [34] and optimistic [5] projections we assume that they implement blocking Gibbs sampling. Since the typical latency of stochastic MTJs stretches from 0.1100.1-10 ns [5, 37], we assume $t_{\text{MTJ}}=2$ ns for the MTJ latency for both the optimisitic and conservative projection.
The main difference between the conservative and optimistic projection is the crossbar latency. The conservative projection uses a crossbar of $256\times 256$ with a latency of $t_{\text{c}}=133$ ns [34] and the optimistic projection uses a crossbar of $100\times 100$ with a latency of $t_{\text{c}}=10$ ps [5].
Every sIM update consists of updating all visible and hidden spins, which requires two crossbar operations, Eq. (24) and (26), and two read operations of the stochastic MTJs, Eq. 25 and 27. For the conservative projection we get,
$\begin{split}T_{\mathrm{sIM}}^{\mathrm{update}}&=2\cdot T_{\text{c}}+2\cdot T_{\text{MTJ}}+T_{\mathrm{acc}}\\ &=2\cdot 133\text{ ns}+2\cdot 2\text{ ns}+130\text{ ns}\approx 400\text{ ns}\end{split}$ (28)
where $T_{\mathrm{sIM}}^{\mathrm{update}}$ is the time per sIM update, $T_{\text{c}}$ is the time for a crossbar operation, $T_{\text{MTJ}}$ is the time for reading the stochastic MTJs, and $T_{\mathrm{acc}}$ is the assumed time to accumulate the results of the up to 64 crossbars on one chip and to round it off to $400$ ns for this rough estimate. For the optimistic projection we get,
$\begin{split}T_{\mathrm{sIM}}^{\mathrm{update}}&=2\cdot T_{\text{c}}+2\cdot T_{\text{MTJ}}+T_{\mathrm{acc}}\\ &=2\cdot 10\text{ ps}+2\cdot 2\text{ ns}\approx 4\text{ ns}\end{split}$ (29)
where $T_{\mathrm{acc}}$ is assumed to be negligible since $T_{\text{c}}\ll T_{\text{MTJ}}$.
Appendix G Comparing energy usage of hardware sIM and MH sampling
The first step is to calculate the energy usage of one sIM update on the hardware of the conservative projection. For $n_{\mathrm{spins}}=484$ and $\alpha=2$, 4 cores of the chip are used, each core uses $100$ mW and $400$ ns per step, resulting in
$E_{\mathrm{sIM}}=4\cdot 100\text{ mW}\cdot 400\text{ ns}=160\text{ nJ}$ (30)
The second step is to estimate the energy usage of one MH sweep on the CPU. We use the reported 180 W TDP of the AMD EPYC 7601 CPU and divide by the 32 cores and the $12$ ms measured average runtime per core while utilizing 16 cores.
$E_{\mathrm{MH}}=\frac{180\text{ W}}{32}\cdot 12\text{ ms}=67.5\text{ mJ}$ (31)
The sampling interval for MH sampling is set to $1$ MH sweep and the sampling interval for sIM sampling is set based on the autocorrelation time ratio using Eq. (14). In this case, the ratio of energy usage for obtaining one sample with MH compared to sIM is given by
$\gamma=\frac{E_{\mathrm{MH}}}{E_{\mathrm{sIM}}}\frac{\Delta t_{\mathrm{MH}}}{\Delta t_{\mathrm{sIM}}f}=\frac{67.5\text{ mJ}}{160\text{ nJ}}\frac{1}{129}\approx 3\cdot 10^{3}$ (32)
where we used that $\Delta t_{\mathrm{sIM}}f=129$ sIM updates for $n=484$ and $\alpha=2$. Therefore, the energy usage of sIM sampling is more than 3 orders of magnitude lower than MH sampling at $n=484$ and $\alpha=2$.
Appendix H Energy barrier analysis
@@
$\Delta E_{i}({z})=H_{\text{ising}}({z}_{i}\rightarrow-{z}_{i})-H_{\text{ising}}({z})$ (33)
where E(z)E({z}) is the energy of the state z{z} given by Eq. (4). The high autocorrelation time for models with higher $\alpha$ is related to the freezing of the visible spins which suggests that there is a high energy barrier for flipping the visible spins. In Fig. 10a, the average measured energy barrier for flipping visible spins of typical samples is shown as a function of n{n} and $\alpha$. The energy barrier can be compared to the logarithm of the autocorrelation time shown in Fig. 10b, which shows qualitatively the same behavior. This suggests that the autocorrelation time is related to the energy barrier of the visible spins. Additionally, the energy barrier of the visible spins can be roughly approximated. For that we first simplify Eq. (33) to
$\Delta E_{i}({z})=2{z}_{i}\sum_{j=1}^{{m}}J_{ij}{z}_{j}+h_{i}{z}_{i}$ (34)
@@
Figure 10: (a) The average energy barrier associated with flipping visible spins of sIM samples, given by $\Delta E=\frac{1}{n}\sum_{i=1}^{n}\frac{1}{{M}}\sum_{\mu=1}^{{M}}\Delta E_{i}({z}^{\mu})=\frac{1}{n{M}}\sum_{\mu=1}^{{M}}\sum_{i=1}^{n}E({z}_{i}^{\mu}\rightarrow-{z}_{i}^{\mu})-E({z}^{\mu})$. (b) The logarithm of the autocorrelation time $\tau_{\mathrm{Evar,sIM}}$. (c) The approximated energy barrier over the weight matrix ${\Delta E^{\mathrm{approx}}_{s_{i}}}=\frac{1}{n}\sum_{i=1}^{n}\sum_{j=1}^{{m}}|W_{ij}|$ as a function of n{n} and $\alpha$. (a, b, c) The models considered have Id = 0 and the results are presented as a function of $\alpha$ and n{n}.
Figure 11: The mean connection strength of the weight matrix, given by $\frac{1}{\alpha n^{2}}\sum_{i,j=1}^{n,{m}}|W_{ij}|$ as a function of n{n} and $\alpha$